What is the correct mapping rule to show the transformation of y=2^x for the equation 1/4(y-5)=2^2x-3 ? Can someone please help, I've been stuck for a few days on this one.I'm guessing: (x,y)-(x+3),(4y+5), anyone know?
Answer by stanbon(75887)

What is the correct mapping rule to show the transformation of y=2^x for the equation 1/4(y-5)=2^2x-3 ? Can someone please help, I've been stuck for a few days on this one.I'm guessing: (x,y)-(x+3),(4y+5), anyone know?
----------
Rearranging I get:
y = 4*(2^(2x-3)-3)+5
y = 4*2^(2x-3)+7
Not sure about this one, but:
It looks to me like y=2^x is moved (3/2)to the right, up 7, and "stretched"
by a factor of 4.
-----------------
Maybe (x,y)----(x-(3/2),4y+7)
----------------
What do you think?
=============
Cheers,
Stan H.